# Bayesian Multiobject Tracking With Neural-Enhanced Motion and Measurement Models

### v2

@@ III-B Data Association and Measurement Model
We consider a conventional MOT model that follows the point object tracking assumption, i.e., each object can generate at most one measurement, and each measurement can originate from at most one object. The former can be encoded by “object-oriented” data association vectors, i.e., $\bm{\mathbf{a}}_{k}=[a_{k}^{1}~{}a_{k}^{2}~{}{\cdots}~{}a_{k}^{I_{k}}]^{\mathrm{T}}$, with $a_{k}^{i}\in\{0,1,\dots,J_{k}\}$. The latter can be encoded by “measurement-oriented” data association vectors, i.e., $\bm{\mathbf{b}}_{k}=[b_{k}^{1}~{}b_{k}^{2}~{}\cdots~{}b_{k}^{J_{k}}]^{\mathrm{T}}$, with $b_{k}^{j}\in\{0,1,\dots,I_{k}\}$. The association vectors have the following relationships: (i) $a_{k}^{i}=j$ indicates that legacy PO $i$ is associated with measurement $j$, while $a_{k}^{i}=0$ indicates that legacy PO $i$ is not associated to any measurement; (ii) $b_{k}^{j}=i$ indicates that measurement $j$ is associated with legacy PO $i$, while $b_{k}^{j}=0$ indicates that measurement $j$ is not associated to any PO. A valid data association event, i.e., an event that satisfies both constraints, can be represented by both object-oriented and measurement-oriented association vectors. On the other hand, if both object-oriented and measurement-oriented association vectors can represent an event, it must be valid.
This insight motivates a hybrid representation of measurement origin uncertainty by both object-oriented and measurement-oriented vectors and to whether $\bm{\mathbf{a}}_{k}$ and $\bm{\mathbf{b}}_{k}$ are consistent [55, 4], i.e.,
@@
Note that the factors in (10) check local pairwise consistency of one entry in $\bm{\mathbf{a}}_{k}$ with one entry in $\bm{\mathbf{b}}_{k}$. If and only if all possible pairs of entries are consistent, we have global consistency, i.e., $\Phi(\bm{\mathbf{a}}_{k},\bm{\mathbf{b}}_{k})=1$ which means that $\bm{\mathbf{a}}_{k}$ and $\bm{\mathbf{b}}_{k}$ represent the same valid data association event [55, 4].
To perform the measurement update step in the presence of measurement-origin uncertainy, we need to compute likelihood ratios for both legacy and new POs. For legacy PO $i$, this likelihood ratio denoted $q(\underline{\bm{\mathbf{x}}}_{k}^{i},\underline{r}_{k}^{i},a_{k}^{i};\bm{\mathbf{z}}_{k})$, is defined as [4]
@@
where $p_{d}$ is the detection probability and $C_{\mathrm{fp}}(\bm{\mathbf{z}}^{j}_{k})=\mu_{\mathrm{fp}}\hskip 0.85358ptf_{\mathrm{fp}}(\bm{\mathbf{z}}^{j}_{k})$ is the intesity of false positive measurements. Furthermore, $1(a_{k}^{i})$ denotes the indicator function, i.e., $11(a_{k}^{i})=$ 1 if $a_{k}^{i}=0$ and 00 otherwise. If a measurement $j\in\{1,\dots,J_{k}\}$ is generated by a PO $i$, it is distributed according to $f(\bm{\mathbf{z}}^{j}_{k}|\underline{\bm{\mathbf{x}}}_{k}^{i})$. This conditional PDF can also be interpreted of affinity of measurement $\bm{\mathbf{z}}^{j}_{k}$ with PO state $\underline{\bm{\mathbf{x}}}_{k}^{i}$. If the measurement $j$ is not generated by any PO, it is a false positive measurement that follows an uninformative PDF denoted $f_{\mathrm{fp}}(\bm{\mathbf{z}}^{j}_{k})$. The number of false positive measurements is modeled by a Poisson PMF with mean $\mu_{\mathrm{fp}}$.
Following the BP framework for MOT [4], the function $l(\bm{\mathbf{z}}_{k}|\underline{\bm{\mathbf{y}}}^{i}_{k})$ used in the measurement update step (1) can be obtained as
$l(\bm{\mathbf{z}}_{k}|\underline{\bm{\mathbf{y}}}^{i}_{k})=\sum^{J_{k}}_{a_{k}^{i}=0}q(\underline{\bm{\mathbf{x}}}_{k}^{i},\underline{r}_{k}^{i},a_{k}^{i};\bm{\mathbf{z}}_{k})\kappa(a_{k}^{i})\vspace{-.5mm}$ (15)
@@
For new PO $j$, the likelihood ratio denoted $v(\overline{\bm{\mathbf{x}}}_{k}^{j},\overline{r}_{k}^{j},b_{k}^{j};\bm{\mathbf{z}}^{j}_{k})$ is defined as [4]
@@
Following the BP framework for MOT [4], the function $n(\overline{\bm{\mathbf{y}}}^{j}_{k};\bm{\mathbf{z}}_{k})$ used in the object initialization step (2), can be obtained as
$n(\overline{\bm{\mathbf{y}}}^{j}_{k};\bm{\mathbf{z}}_{k})=\sum^{I_{k}}_{b_{k}^{j}=0}v(\overline{\bm{\mathbf{x}}}_{k}^{j},\overline{r}_{k}^{j},b_{k}^{j};\bm{\mathbf{z}}^{j}_{k})\iota(b_{k}^{j})\vspace{-.5mm}$ (20)
where the data association functions $\iota(b_{k}^{j})$ are provided loopy BP [4], making use of the hybrid data association representation and the corresponding consistency constraint (9).
A detailed derivation of likelihood ratios $q(\underline{\bm{\mathbf{x}}}_{k}^{i},\underline{r}_{k}^{i},a_{k}^{i};\bm{\mathbf{z}}_{k})$ and $v(\overline{\bm{\mathbf{x}}}_{k}^{j},\overline{r}_{k}^{j},b_{k}^{j};\bm{\mathbf{z}}^{j}_{k})$ and the corresponding measurement update steps is also provided in [4]. In Section V, we will discuss how the likelihood ratios can be neural enhanced by incorporating information learned from raw data not subject to measurement detection.
@@ III-C Object Declaration and State Estimation
In our Bayesian setting, we declare the existence of POs and estimate their states based on the marginal existence probabilities $p(r^{i}_{k}=1|\bm{\mathbf{z}}_{1:k})$ and the conditional PDF $f(\bm{\mathbf{x}}^{i}_{k}|r^{i}_{k}=1,\bm{\mathbf{z}}_{1:k})$. More specifically, a PO is declared to exist if $p(r^{i}_{k}=1|\bm{\mathbf{z}}_{1:k})$ is above a predefined threshold $T_{\mathrm{dec}}$, e.g., $T_{\mathrm{dec}}=0.5$. For PO $i$ that is declared to exist, a state estimate of $\bm{\mathbf{x}}^{i}_{k}$ is then provided by the minimum-mean-square-error (MMSE) estimator [29], i.e.,
$\hat{\bm{\mathbf{x}}}^{i}_{k}=\int\bm{\mathbf{x}}^{i}_{k}\hskip 0.85358ptf(\bm{\mathbf{x}}^{i}_{k}\hskip 0.85358pt|\hskip 0.85358ptr^{i}_{k}=1,\bm{\mathbf{z}}_{1:k})\hskip 0.85358pt\mathrm{d}\bm{\mathbf{x}}^{i}_{k}.$ (21)
Note that the existence probability and the conditional PDF are computed as $p(r^{i}_{k}=1\hskip 0.85358pt|\hskip 0.85358pt\bm{\mathbf{z}}_{1:k})=\int f(\bm{\mathbf{x}}^{i}_{k},r^{i}_{k}=1\hskip 0.85358pt|$ $\bm{\mathbf{z}}_{1:k})\mathrm{d}\bm{\mathbf{x}}^{i}_{k}$ and $f(\bm{\mathbf{x}}^{i}_{k}\hskip 0.85358pt|\hskip 0.85358ptr^{i}_{k}=1,\bm{\mathbf{z}}_{1:k})=f(\bm{\mathbf{y}}^{i}_{k}\hskip 0.85358pt|\hskip 0.85358pt\bm{\mathbf{z}}_{1:k})/p(r^{i}_{k}=1\hskip 0.85358pt|\hskip 0.85358pt\bm{\mathbf{z}}_{1:k})$.
As discussed previously, in this work we use belief propagation [4] to compute the marginal posterior PDFs $f(\bm{\mathbf{y}}^{i}_{k}\hskip 0.85358pt|\hskip 0.85358pt\bm{\mathbf{z}}_{1:k})=f(\bm{\mathbf{x}}^{i}_{k},r^{i}_{k}\hskip 0.85358pt|\hskip 0.85358pt\bm{\mathbf{z}}_{1:k})$. Note that, since a new PO is introduced for each measurement, the number of POs grows with time $k$. Thus, to limit computational complexity, POs whose approximate existence probability is below a threshold $T_{\mathrm{pru}}$ are removed from the state space.
@@ IV Prediction Step With Neural Enhanced Motion Model
In Bayesian MOT, the prediction step propagates the posterior PDF from the previous time step of an object’s state through the motion model to obtain a predicted PDF at the current time step. In this section, we develop a non-Markovian learnable motion model suitable for neural enhancement, introduce the corresponding neural architecture, and discuss how the resulting prediction step can be implemented using SPs.
IV-A Non-Markovian, Learnable Motion Model
@@
We assume that conditioned on ${\bm{\mathbf{x}}}^{i}_{k_{i}:k-1}$, and ${\mathbf{s}}^{i}_{k_{i}:k-1}$, the current kinematic state ${\underline{\bm{\mathbf{x}}}}^{i}_{k}$ is independent of all previous states of other objects that are not in its neighborhood or do not exisit, i.e.,
$f({\underline{\bm{\mathbf{x}}}}^{i}_{k}|{\bm{\mathbf{x}}}^{i}_{k_{i}:k-1},{\mathbf{y}}^{\sim i}_{k_{i}:k-1})=f({\underline{\bm{\mathbf{x}}}}^{i}_{k}|{\bm{\mathbf{x}}}^{i}_{k_{i}:k-1},{\mathbf{s}}^{i}_{k_{i}:k-1}).$ (22)
In addition, it is assumed that there is a high-dimensional hidden state $\bm{\mathbf{h}}^{i}_{k-1}\in\mathbb{R}^{d_{\mathrm{h}}}$ that is given by $\bm{\mathbf{h}}^{i}_{k-1}=T\big{(}{\bm{\mathbf{x}}}^{i}_{k_{i}:k-2},{\mathbf{s}}^{i}_{k_{i}:k-2}\big{)}$ such that
@@
Conditioned on ${\bm{\mathbf{x}}}^{i}_{k-1}$ and ${\mathbf{s}}^{i}_{k-1}$, the hidden state is a sufficient statistic of ${\bm{\mathbf{x}}}^{i}_{k_{i}:k-2}$ and ${\mathbf{s}}^{i}_{k_{i}:k-2}$ with respect to ${\underline{\bm{\mathbf{x}}}}^{i}_{k}$. Based the assumptions discussed above, the transition function in (4) can be denoted as
$f(\underline{\bm{\mathbf{y}}}^{i}_{k}|{\bm{\mathbf{y}}}_{k_{i}:k-1})=f(\underline{\bm{\mathbf{x}}}^{i}_{k},\underline{r}^{i}_{k}|\bm{\mathbf{x}}^{i}_{k-1},r^{i}_{k-1},{\mathbf{s}}^{i}_{k-1},\bm{\mathbf{h}}^{i}_{k-1}).$ (24)
@@
The motion model related to $f({\underline{\bm{\mathbf{x}}}}^{i}_{k}|{\bm{\mathbf{x}}}^{i}_{k-1},{\mathbf{s}}^{i}_{k-1},\bm{\mathbf{h}}^{i}_{k-1})$ in (23) is assumed as
${\underline{\bm{\mathbf{x}}}}^{i}_{k}=g_{1}(\bm{\mathbf{x}}^{i}_{k-1},{\mathbf{s}}^{i}_{k-1},\bm{\mathbf{h}}^{i}_{k-1})+\bm{\mathbf{v}}^{i}_{k}$ (25)
@@
For future reference, we introduce the joint motion model $g(\bm{\mathbf{x}}^{i}_{k-1},{\mathbf{s}}^{i}_{k-1},\bm{\mathbf{h}}^{i}_{k-1})\triangleq\big{[}g_{1}(\bm{\mathbf{x}}^{i}_{k-1},{\mathbf{s}}^{i}_{k-1},\bm{\mathbf{h}}^{i}_{k-1})^{\mathrm{T}}\hskip 0.85358pt\hskip 0.85358ptg_{2}(\bm{\mathbf{x}}^{i}_{k-1},$ $\bm{\mathbf{s}}^{i}_{k-1},\bm{\mathbf{h}}^{i}_{k-1})^{\mathrm{T}}\big{]}^{\mathrm{T}}$.

### v3

@@ III-B Data Association and Measurement Model
We consider a conventional MOT model that follows the point object tracking assumption, i.e., each object can generate at most one measurement, and each measurement can originate from at most one object. The former can be encoded by “object-oriented” data association vectors, i.e., $\bm{\mathbf{a}}_{k}=[a_{k}^{1}~a_{k}^{2}~{\cdots}~a_{k}^{I_{k}}]^{\mathrm{T}}$, with $a_{k}^{i}\in\{0,1,\dots,J_{k}\}$. The latter can be encoded by “measurement-oriented” data association vectors, i.e., $\bm{\mathbf{b}}_{k}=[b_{k}^{1}~b_{k}^{2}~\cdots~b_{k}^{J_{k}}]^{\mathrm{T}}$, with $b_{k}^{j}\in\{0,1,\dots,I_{k}\}$. The association vectors have the following relationships: (i) $a_{k}^{i}\hskip-0.85358pt=\hskip-0.85358ptj$ indicates that legacy PO $i$ is associated with measurement $j$, while $a_{k}^{i}=0$ indicates that legacy PO $i$ is not associated to any measurement; (ii) $b_{k}^{j}=i$ indicates that measurement $j$ is associated with legacy PO $i$, while $b_{k}^{j}=0$ indicates that measurement $j$ is not associated to any PO. A valid data association event, i.e., an event that satisfies both constraints, can be represented by both object-oriented and measurement-oriented association vectors. On the other hand, if both object-oriented and measurement-oriented association vectors can represent an event, it must be valid.
This insight motivates a hybrid representation of measurement origin uncertainty by both object-oriented and measurement-oriented vectors and to check whether $\bm{\mathbf{a}}_{k}$ and $\bm{\mathbf{b}}_{k}$ are consistent [66, 38], i.e.
@@
Note that the factors in (9) check local pairwise consistency of one entry in $\bm{\mathbf{a}}_{k}$ with one entry in $\bm{\mathbf{b}}_{k}$. If and only if all possible pairs of entries are consistent, we have global consistency, i.e., $\Phi(\bm{\mathbf{a}}_{k},\bm{\mathbf{b}}_{k})\hskip-0.85358pt=\hskip-0.85358pt1$ which means that $\bm{\mathbf{a}}_{k}$ and $\bm{\mathbf{b}}_{k}$ represent the same valid data association event [66, 38].
To perform the measurement update step in the presence of measurement-origin uncertainty, we need to compute likelihood ratios for both legacy and new POs. For legacy PO $i$, this likelihood ratio denoted $q(\underline{\bm{\mathbf{x}}}_{k}^{i},\underline{r}_{k}^{i},a_{k}^{i};\bm{\mathbf{z}}_{k})$, is defined
as [38]
@@
where $p_{d}$ is the detection probability and $C_{\mathrm{fp}}(\bm{\mathbf{z}}^{j}_{k})=\mu_{\mathrm{fp}}\hskip 0.85358ptf_{\mathrm{fp}}(\bm{\mathbf{z}}^{j}_{k})$ is the intensity of false positive measurements. Furthermore, $1(a_{k}^{i})$ denotes the indicator function, i.e., $1(a_{k}^{i})=1$ if $a_{k}^{i}=0$ and 0 otherwise. If a measurement $j\in\{1,\dots,J_{k}\}$ is generated by a PO $i$, it is distributed according to $f(\bm{\mathbf{z}}^{j}_{k}|\underline{\bm{\mathbf{x}}}_{k}^{i})$. This conditional PDF can also be interpreted of affinity of measurement $\bm{\mathbf{z}}^{j}_{k}$ with PO state $\underline{\bm{\mathbf{x}}}_{k}^{i}$. If the measurement $j$ is not generated by any PO, it is a false positive measurement that follows an uninformative PDF denoted $f_{\mathrm{fp}}(\bm{\mathbf{z}}^{j}_{k})$. The number of false positive measurements is modeled by a Poisson PMF with mean $\mu_{\mathrm{fp}}$.
Following the BP framework for MOT [38], the function $l(\bm{\mathbf{z}}_{k}|\underline{\bm{\mathbf{y}}}^{i}_{k})$ used in the measurement update step (1) can be obtained
as
$l(\bm{\mathbf{z}}_{k}|\underline{\bm{\mathbf{y}}}^{i}_{k})=\sum^{J_{k}}_{a_{k}^{i}=0}q(\underline{\bm{\mathbf{x}}}_{k}^{i},\underline{r}_{k}^{i},a_{k}^{i};\bm{\mathbf{z}}_{k})\kappa(a_{k}^{i})\vskip-1.42262pt$ (14)
@@
For new PO $j$, the likelihood ratio denoted $v(\overline{\bm{\mathbf{x}}}_{k}^{j},\overline{r}_{k}^{j},b_{k}^{j};\bm{\mathbf{z}}^{j}_{k})$ is defined
as [38]
@@
Following the BP framework for MOT [38], the function $n(\overline{\bm{\mathbf{y}}}^{j}_{k};\bm{\mathbf{z}}_{k})$ used in the object initialization step (2), can be obtained
as
$n(\overline{\bm{\mathbf{y}}}^{j}_{k};\bm{\mathbf{z}}_{k})=\sum^{I_{k}}_{b_{k}^{j}=0}v(\overline{\bm{\mathbf{x}}}_{k}^{j},\overline{r}_{k}^{j},b_{k}^{j};\bm{\mathbf{z}}^{j}_{k})\iota(b_{k}^{j})\vskip-1.42262pt$ (19)
where the data association functions $\iota(b_{k}^{j})$ are provided by loopy BP [38], making use of the hybrid data association representation and the corresponding consistency constraint (8).
A detailed derivation of likelihood ratios $q(\underline{\bm{\mathbf{x}}}_{k}^{i},\underline{r}_{k}^{i},a_{k}^{i};\bm{\mathbf{z}}_{k})$ and $v(\overline{\bm{\mathbf{x}}}_{k}^{j},\overline{r}_{k}^{j},b_{k}^{j};\bm{\mathbf{z}}^{j}_{k})$ and the corresponding measurement update steps is also provided in [38]. In Section V, we will discuss how the likelihood ratios can be neural enhanced by incorporating information learned from raw data not subject to measurement detection
@@ III-C Object Declaration and State Estimation
In our Bayesian setting, we declare the existence of POs and estimate their states based on the marginal existence probabilities $p(r^{i}_{k}\hskip-0.85358pt=\hskip-0.85358pt1|\bm{\mathbf{z}}_{1:k})$ and the conditional PDF $f(\bm{\mathbf{x}}^{i}_{k}|r^{i}_{k}\hskip-0.85358pt\hskip-0.85358pt=1,\bm{\mathbf{z}}_{1:k})$. More specifically, a PO is declared to exist if $p(r^{i}_{k}\hskip-0.85358pt=\hskip-0.85358pt1|\bm{\mathbf{z}}_{1:k})$ is above a predefined threshold $T_{\mathrm{dec}}$, e.g., $T_{\mathrm{dec}}=0.5$. For PO $i$ that is declared to exist, a state estimate of $\bm{\mathbf{x}}^{i}_{k}$ is then provided by the MMSE estimator [21], i.e.
$\hat{\bm{\mathbf{x}}}^{i}_{k}\hskip-0.85358pt=\hskip-0.85358pt\int\bm{\mathbf{x}}^{i}_{k}\hskip 0.85358ptf(\bm{\mathbf{x}}^{i}_{k}\hskip 0.85358pt|\hskip 0.85358ptr^{i}_{k}\hskip-0.85358pt=\hskip-0.85358pt1,\bm{\mathbf{z}}_{1:k})\hskip 0.85358pt\mathrm{d}\bm{\mathbf{x}}^{i}_{k}.$
Note that the existence probability and the conditional PDF are computed as $p(r^{i}_{k}\hskip-0.85358pt=1\hskip 0.85358pt|\hskip 0.85358pt\bm{\mathbf{z}}_{1:k})\hskip-0.85358pt=\hskip-0.85358pt\int f(\bm{\mathbf{x}}^{i}_{k},r^{i}_{k}\hskip-0.85358pt=\hskip-0.85358pt1\hskip 0.85358pt|$ $\bm{\mathbf{z}}_{1:k})\mathrm{d}\bm{\mathbf{x}}^{i}_{k}$ and $f(\bm{\mathbf{x}}^{i}_{k}\hskip 0.85358pt|\hskip 0.85358ptr^{i}_{k}\hskip-0.85358pt=\hskip-0.85358pt1,\bm{\mathbf{z}}_{1:k})\hskip-0.85358pt=\hskip-0.85358ptf(\bm{\mathbf{y}}^{i}_{k}\hskip 0.85358pt|\hskip 0.85358pt\bm{\mathbf{z}}_{1:k})/p(r^{i}_{k}=1\hskip 0.85358pt|\hskip 0.85358pt\bm{\mathbf{z}}_{1:k})$.
As discussed previously, in this work we use BP [38] to compute the marginal posterior PDFs $f(\bm{\mathbf{y}}^{i}_{k}\hskip 0.85358pt|\hskip 0.85358pt\bm{\mathbf{z}}_{1:k})\hskip-0.85358pt=f(\bm{\mathbf{x}}^{i}_{k},r^{i}_{k}\hskip 0.85358pt|\hskip 0.85358pt\bm{\mathbf{z}}_{1:k})$. Note that, since a new PO is introduced for each measurement, the number of POs grows with time $k$. Thus, to limit computational complexity, POs whose approximate existence probability is below a threshold $T_{\mathrm{pru}}$ are removed from the state space
@@ IV Prediction Step With Neural Enhanced Motion Model
The goal of this section is to compute a predicted PDF that takes interactions with previous object states and neighboring object states into account. This requires addressing three challenges: (1) developing a non-Markovian motion model that involves neighboring object states, (2) establishing a computational framework to implement the motion model, and (3) performing a high-dimensional marginalization integral to compute the predicted PDF. To address these challenges, we design a motion model that can be naturally implemented by neural networks, and discuss how the predicted PDF is computed using SPs.
IV-A Non-Markovian Motion Model With Object Interactions
@@
We assume that conditioned on ${\bm{\mathbf{x}}}^{i}_{k_{i}:k-1}$, and ${\mathbf{s}}^{i}_{k_{i}:k-1}$, the current kinematic state ${\underline{\bm{\mathbf{x}}}}^{i}_{k}$ is independent of all previous states of other objects that are not in its neighborhood or do not exist, i.e.
$f({\underline{\bm{\mathbf{x}}}}^{i}_{k}|{\bm{\mathbf{x}}}^{i}_{k_{i}:k-1},{\mathbf{y}}^{\sim i}_{k_{i}:k-1})=f({\underline{\bm{\mathbf{x}}}}^{i}_{k}|{\bm{\mathbf{x}}}^{i}_{k_{i}:k-1},{\mathbf{s}}^{i}_{k_{i}:k-1}).$
In addition, it is assumed that there is a high-dimensional hidden state $\bm{\mathbf{h}}^{i}_{k-1}\hskip-0.85358pt\hskip-0.85358pt\in\hskip-0.85358pt\hskip-0.85358pt\mathbb{R}^{d_{\mathrm{h}}}$ that is given by $\bm{\mathbf{h}}^{i}_{k-1}=T\big({\bm{\mathbf{x}}}^{i}_{k_{i}:k-2},{\mathbf{s}}^{i}_{k_{i}:k-2}\big)$ such that
@@
Conditioned on ${\bm{\mathbf{x}}}^{i}_{k-1}$ and ${\mathbf{s}}^{i}_{k-1}$, the hidden state is a sufficient statistic of ${\bm{\mathbf{x}}}^{i}_{k_{i}:k-2}$ and ${\mathbf{s}}^{i}_{k_{i}:k-2}$ with respect to ${\underline{\bm{\mathbf{x}}}}^{i}_{k}$. Based these assumptions, the transition function in (3) can be denoted as
$f(\underline{\bm{\mathbf{y}}}^{i}_{k}|{\bm{\mathbf{y}}}_{k_{i}:k-1})=f(\underline{\bm{\mathbf{x}}}^{i}_{k},\underline{r}^{i}_{k}|\bm{\mathbf{x}}^{i}_{k-1},r^{i}_{k-1},{\mathbf{s}}^{i}_{k-1},\bm{\mathbf{h}}^{i}_{k-1}).$
@@
The motion model related to $f({\underline{\bm{\mathbf{x}}}}^{i}_{k}|{\bm{\mathbf{x}}}^{i}_{k-1},{\mathbf{s}}^{i}_{k-1},\bm{\mathbf{h}}^{i}_{k-1})$ in (20) is assumed
as
${\underline{\bm{\mathbf{x}}}}^{i}_{k}=g_{1}(\bm{\mathbf{x}}^{i}_{k-1},{\mathbf{s}}^{i}_{k-1},\bm{\mathbf{h}}^{i}_{k-1})+\bm{\mathbf{v}}^{i}_{k}$
@@
For future reference, we introduce the joint motion model $g(\bm{\mathbf{x}}^{i}_{k-1},{\mathbf{s}}^{i}_{k-1},\bm{\mathbf{h}}^{i}_{k-1})\triangleq\big[g_{1}(\bm{\mathbf{x}}^{i}_{k-1},{\mathbf{s}}^{i}_{k-1},\bm{\mathbf{h}}^{i}_{k-1})^{\mathrm{T}}\hskip 0.85358pt\hskip 0.85358ptg_{2}(\bm{\mathbf{x}}^{i}_{k-1},$ $\bm{\mathbf{s}}^{i}_{k-1},\bm{\mathbf{h}}^{i}_{k-1})^{\mathrm{T}}\big]^{\mathrm{T}}\hskip-0.85358pt\hskip-0.85358pt$.